OpenStudy (anonymous):

The perimeters of two similar polygons have the same ratio as any two corresponding sides. This can be expressed by the formula..

OpenStudy (anonymous):

The perimeter of two similar polygons have the same ratio as any two corresponding sides. This can be expressed by the formula \[P/ P^{'}=a/a ^{'}\] Find P' when P= 210 yards, a=2. and a'=3